OpenStudy (anonymous):

A heat engine has a cold reservoir of 250°C and a hot reservoir of 715°C. What is the efficiency of the engine?

OpenStudy (sidsiddhartha):

max efficiency of a heat engine is\[\frac{ (T1-T2) }{ T1 }=1-\frac{ T2 }{ T1 }\] ur T1=715c=(715+273)=988k and T2=250c=23k now just plug in those values :)